A company issues 7%, 10-year bonds with a face amount of $80,000 for $74,564 on January 1, 2021. The market interest rate for bonds of similar risk and maturity is 8%. Interest is paid semiannually on June 30 and December 31.

Required:
1. & 2. Record the bond issue and first interest payment on June 30, 2021. (If no entry is required for a particular transaction/event, select "No Journal Entry Required" in the first account field. Do not round intermediate calculations. Round your answers to the nearest dollar amount.)

Journal entry worksheet

  • Record the bond issue.

Note: Enter debits before credits.

Date General Journal Debit Credit
January 01, 2021

Journal entry worksheet

  • Record the first interest payment on June 30, 2021.

Note: Enter debits before credits.

Date General Journal Debit Credit
June 30, 2021
Answer #1
Debit Credit
January 01,2021 Cash 74564
Discount on Bonds payable 5436
     Bonds payable 80000
Debit Credit
June 30,2021 Interest expense 2983 =74564*8%/2
      Discount on Bonds payable 183
      Cash 2800 =80000*7%/2
Alternatively, if straight line method is used:
Debit Credit
June 30,2021 Interest expense 3072
      Discount on Bonds payable 272 =5436/20
      Cash 2800 =80000*7%/2